se the following table to answer the next question. The base year is 2007.

Year

Hot Dogs

Price           Quantity

Baseballs

Price           Quantity

Bottles of Beer

Price          Quantity

2005

$ 2.50             100

$ 2.50                  50

$1.00               100

2006

    4.00             100

    5.00                100

2.00               150

2007

    5.00             100

    5.00                100

2.00               200

2008

    8.00             150

    8.00                200

4.00               200

2009

10.00             200

10.00                200

4.00               250

The Consumer Price Index for the year 2005 is
Answer #1

Answer
Cost of a basket in base year =sum of (the price of the base year * quantity of base year)
Cost of the basket in base year=5*100+5*100+2*200
=1400
Cost of a basket in current year =sum of (the price of the current year* quantity of base year)
Cost of basket in 2005=2.5*100+2.5*100+1*200
=700
CPI =(Cost of the basket in the year/cost of the basket in the base year)*100
=(700/1400)*100
=50

the CPI in the year 2005 is 50